Daily Reports

Because of the pandemic, no observers can come to the telescope. This pool is entirely in remote mode.

Dec 9th, Wednesday

16h - 21h; Bilal: Difficulties to start and get a good focus. Increasingly bad opacities.

21h - 00h; Emmanuel, Bilal & Laurence: Opacity rising from 0.5 to above 1. We observed 126-20 until around 00:30 UT. At 00:45 UT the telescope was stopped due to heavy humidity level.

Dec 10th, Thursday

00h - 03h; Rain

03h - 09h; Rain

09h - 15h; Rain

15h - 21h; Rain

21h - 00h; Rain

Dec 11th, Friday

00h - 03h; Rain

03h - 09h; Rain

09h - 15h; Rain

15h - 21h; Rain

21h - 00h; Rain

Dec 12th, Saturday

00h - 03h; Wind

03h - 09h; Wind

09h - 15h; Wind

15h - 21h; Bilal: Better. We start operations. The opacity is around 0.4. We start the due calibrations (Uranus, and secondary calibrators)

21h - 00h; Marco and Antonio:

Dec 13th, Sunday

00h - 03h; Marco and Antonio:

03h - 09h; Nicolas: Opacity is ok. Continuing on project 096-20 and continuing project 182-20

09h - 15h; Jean-Fran├žois: Opacity is good, we continue project 182-20 and start project 122-16 on the source L1689B. It's extremely close to the sun. Moving to project 126-20.

15h - 21h; Bilal: Continuing project 122-16, and moving to project 126-20 as beam is very distorted and as opacity increases. Once the sun sets, we'll move back to project 134-20.

21h - 00h; Juan, Bilal & Miren: a skydip and 2 repetitions on IC443G (134-20). After another skydip, with tau225 around 0.5, changed to 126-20.

Dec 14th, Monday

00h - 03h; Juan, Bilal & Miren: Continuing on projet 126-20

03h - 09h; Matthieu: Continuing on projet 126-20

09h - 15h; Jean-Fran├žois: Tau is around 0.6-0.7. Continuing on project 126-20.

15h - 21h; Bilal: continuing on projet 126-20. Opacity should decrease the evening.

21h - 00h; Juan & Emmanuel: The opacity is again is better range, we start projet 138-20.

Dec 15th, Tuesday

00h - 03h; Fog

03h - 10h; Fog

DailyReportsNika2Pool091220 (last edited 2020-12-15 14:03:38 by NikaBolometer)